RC Update and Run plan for Sep 19,2020

From HallCWiki
Jump to: navigation, search

RC Meeting Information

RC Daily Meetings are M-F at 4:15 PM

To join the Meeting:  https://bluejeans.com/861439920
To join via phone :  +1.408.740.7256 (US (San Jose)), +1.888.240.2560 (US Toll Free))

Saturday 19 Sept 2020.

General updates

(See d2n Run Plan 2020 for broader status/overview.)

  • Beam energy measurement started at Fri 2:15pm [1], completed around Fri 5:30pm (including restoring production optics)
    • Results: 2182.708 +/- 2.3662 MeV [2], [3]
  • Fri 5:15pm--7:45pm Ion chamber calibrations on No Target, Carbon-Optics, and Ref Cell(135psig N2):* Will continue with 1-pass runplan
  • Fri 7:55pm Began 1-pass production
  • Fri 11:30pm SHMS dipole tripped, not clear why.. Reset cleanly though [4]
    • Run plan delivered to shift crew: 1-pass running program
    • -Done- Beam Energy Measurement (1 hour)
    • -Done- Reference Cell Empty, Hydrogen (H2), Carbon Cross Sections (3--4 hours)
    • -Done- No Target Run
    • -Done- 3He Reference Cell Runs (N2 and 3He Pressure Curves, False Asymmetries) (5hrs ideal, 3 hrs min)
    • -Done- Reference Cell He3 Cross Sections (1 hour)
    • -- In Progress -- Polarized 3He Runs, 3He elastic and QE longitudinal Asymmetries (16 hours)
      • ~50% complete as of Sat, 2pm
      • Flipped Target polarization to Longitudinal/0 around Sat, 2:30pm [5]
      • Another ~8 hours of production follows
    • Polarized 3He Elastic Longitudinal Asymmetry (until end of 1-pass running)
    • NOTE: OPS/MCC would like to stop beam at 7am for some Accel tests
  • Some beam energy stability issues have been noticed by Hall B (in particular); Jay is looking into this
    • These should not impact Hall C significantly, but may feed into beam polarization corrections.
    • Jay has link some of this discussion to the Hall C beam energy measurement HCLog entry and cross posted to the HCLog. This will help us follow breadcrumbs during beam polarization analysis for 1-pass.
    • Hall C beam energy drifted up after the Hall C beam energy measurement was taken [6]
    • Hall B Moller Fri, 09/18/2020 - 10:20 was higher than expected: -86.233 +/- 1.494 (several percent) [7]
    • Energy shift that would impact Hall B polarization at few percent level [8]
    • Energy lock switched from 2A to 3C Sat (19th) @ 9:40 [9]. No notice to Halls, unfortunately..


Notes for the Future

  • 1-pass running program
  • MCC/OPs is requesting that beam delivery to Halls be halted at 7am instead of 8am to run a couple measurements
    • Measurements are intended to guide future work on beam and energy stability issues noted earlier.
  • Hall will move to Controlled Access on Monday
    • When is still TBD; depends when RadCon is available
    • Could be in the afternoon
    • EPR photodiode will be replaced on Monday (JP, Junhao, RadCon)
    • Will move to Restricted Access ________ ?
  • Bill, Murchhana check compass/field measurements
    • Also checking with Jixie re: modeled fields
    • Need to make a write-up that clearly defines all the different coordinate systems used, swapped power supply/polarity issues, etc.

Pending Issues

Opportunistic Jobs/Access

Current Run Plan

Friday SWING → Monday

  • See task list at Run Plan for 1-pass for more details
    • Priority will be PbPt polarization cross check with longitudinal pol. elastics.
    • These have lower priority but will likely be pursued first while target spins up
      • Beam energy measurement (1st thing)
      • 3He, N2 density measurements
      • H2 xsec measurement
      • Pressure curves
  • NOTE: Hall C is in control of the Charge Asymmetry feedback system until Monday
    • Charge Feedback Parity DAQ Instructions
    • We will have to run the ./FlipPCvaluesHallC script in the Parity DAQ VNC session after every IHWP change


Standing Orders

  • Change the IHWP at least every 8 hours when running on glass cells. Call Halls A and B in advance. If there are any issues, call the RC
    • NOTE: Hall C is in control of the Charge Asymmetry feedback system until Monday Charge Feedback Parity DAQ
    • We will have to run the ./FlipPCvaluesHallC script in the Parity DAQ VNC session after every IHWP change
  • Ensure that all information in the CODA run-start GUIs is correct for every run!
  • If beam goes away for more than 3h, or MCC does operations on the beam that may change the beam tune, then move target to No Target position,
    • verify Harps scans at 3H07A and 3H07B are within spec, and
    • ensure beam is stable (with all locks ON) at 0, 0mm before continuing.
    • Beam recovery procedure has more detailed procedure if needed.
  • New beam positions have been verified and established to be (0,0) mm for all targets.
    • Monitor that beam is at (0,0) using the stripchart on the Wall screen and verify no scraping on the glass cells using the Analysis monitor.
  • Check the position locks are on.
  • Check that CODA is running normally
  • Check online replays shortly after starting runs, & full replay results too.
  • Fill BTA with realistic numbers. You may follow the guidance coming from "load from EPICS", if it does not overestimate the ABU (as it generally does). https://logbooks.jlab.org/entry/3783171. If we have to lower the beam current we actually want because the beam is too unstable, prorate the ABUs.
!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!
!!                                                     !!
!!  ATTN: SHIFT LEADERS                                !!
!!     Please follow the shift summary template on     !!
!!     the laminated SL instructions.                  !!
!!                                                     !!
!!  ATTN: TARGET OPERATORS                             !!
!!  All NMR measurement HCLOG entries should include:  !!
!!     1) A title that clearly lists polarization      !!
!!     2) Screenshots of both NMR_FieldSweep and       !!
!!        the polarization fit                         !!
!!                                                     !!
!!     For an example see the following entry:         !!
!!     https://logbooks.jlab.org/entry/3768696       !!
!!                                                     !!
!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!

Notices in Effect for Shift Crew

  • Policy for CANS-related hall status change: Make a logbook entry when the hall goes in, or comes out of laser-controlled access.
  • Magnet policies
    • Follow the updated guidelines in the How-to#Spectrometer_Magnets Spectrometer Magnets How to
    • Spectrometer GUI (go_magnets) should always be visible on Shift Leader computer, hcdesk3
    • If any magnet trips, snapshot the associated power supply window and post it to hclog before resetting the interlock.
  • Monitor the beam position on target and call MCC immediately if you see large excursions (1 mm level).
  • When rotating the spectrometer, use HCCAM01 to keep an eye on the SOLID Cherenkov test stand when rotating SHMS (in addition to watching the live angle-cam image, etc). The updated rotation instructions are here: https://hallcweb.jlab.org/wiki/index.php/Spectrometer_Rotation#NOTICES_IN_EFFECT

Useful Links

PolHe3 Polarization Measurements

Misc Links